Find the percentile rank of the 18 using the following data: 15, 13, 6, 5, 12, 22, 50, 18

1
Expert's answer
2022-05-02T16:30:44-0400

In order to compute the corresponding percentile rank for "X = 18," the data needs to be organized into ascending order, as shown in the table below


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Position & X (Asc. Order) \\ \hline
 1 & 5 \\
 \hdashline
 2 & 6 \\
 \hdashline
 3 & 12 \\
 \hdashline
 4 & 13 \\
 \hdashline
5 & 15 \\
 \hdashline
 6 & 18 \\
 \hdashline
 7 & 22 \\
 \hdashline
8 & 50 \\
 \hdashline
\end{array}"

Let "k" be the number of values in the dataset that are less than "X = 18."

In this case, "k = 5." Hence, the percentile rank is:


"Percentile \ Rank=\dfrac{k}{n}=\dfrac{5}{8}=0.625"

Percentile Rank "=62.5\%."